# Spin transport and lack of quantisation in the $A\mathrm{II}$ class on the honeycomb structure

### v1

@@ 4.3. Lack of quantisation in the $\mathrm{AII}^{\hexagon}$ class
Note that for $m\neq 0$ the integrand in (4.24) is continuous in $\lambda_{\mathrm{R}}$, in a neighbourhood of 00, and uniformly bounded by some integrable function (decaying in $k_{0}$ and constant in $k_{1}$ and $k_{2}$). Thus, by the dominated convergence theorem $\sigma^{s}_{12}$ is also continuous in $\lambda_{\mathrm{R}}$, in a neighbourhood of 00.
@@ A.1. Grand canonical formalism
$\{a(f),a(g)\}=\{a^{*}(f),a^{*}(g)\}=0\;,\qquad\{a^{*}(f),a(g)\}=\langle f,g\rangle_{\mathcal{H}}\;,\qquad\forall f,g\in\mathcal{H}\;,$
@@
$\alpha_{t}\big{(}a(f)\big{)}:=a(\mathrm{e}^{\mathrm{i}t(H-\mu\mathbb{1})}f)\;,\qquad\forall f\in\mathcal{H}\;.$ (A.3)
More precisely, one introduces the set of states that satisfy the KMS property with respect to $(\alpha_{t})_{t\in\mathbb{R}}$, see [BR, Definition 5.3.1].
@@
Following the proof of Lemma A.10, one can see that the second and third terms are $O(L^{2-1/\xi})$, $\xi$ being the constant associated with $B$ according to (2.5), so that the first claim is proven. Concerning (A.10), since by Lemma 3.5 $[B,X]\in\mathcal{P}(\mathcal{H})$ we can apply Lemma A.10 and use that $\chi_{L}$ commutes with $X$. ∎
@@
$[\mathrm{d}\Gamma(A_{L}),\mathrm{d}\Gamma(B)]=\mathrm{d}\Gamma([A_{L},B])=\lim_{R\to\infty}\mathrm{d}\Gamma([A_{L},B_{R}])$
@@
since $\gamma^{H,\mu}_{\beta}$ is bounded. Then, we can take the $\beta\to\infty$ as in (A.12) since $[A_{L},B]$ is trace class, and conclude by (A.9).
Finally, to prove the last claim we use that $\chi_{L}$ is a projection and that it commutes with $X$, so that $[B_{L},X]=[B_{L},X_{L}]$. But $X_{L}$ is bounded hence by (A.6)
@@
where we also used that $X=X_{L}+\chi^{c}_{L}X$ and that $[\mathrm{d}\Gamma(B_{L}),\mathrm{d}\Gamma(\chi^{c}_{L}X)]=\mathrm{d}\Gamma\big{(}[B_{L},\chi^{c}_{L}X]\big{)}=0$. The conclusion follows by arguing as before, and by using (A.10). ∎
@@ A.2. Wick’s rotation of the Kubo formula
To begin, we provide a Kubo formula for the spin conductivity in $\mathrm{AII}^{\hexagon}$. We henceforth denote $J:=\mathrm{i}[H,X]$ the charge current operator associated with the Hamiltonian $H$.
@@
Let $(H,\mu)$ be an insulator with $H\in\mathrm{AII}^{\hexagon}$ and let $P=P(H,\mu)$ be the associated Fermi projector. Let $J_{i}^{\mathrm{conv}}$ be the conventional spin current and let $J(s):=\mathrm{e}^{\mathrm{i}sH}J\mathrm{e}^{-\mathrm{i}sH}$ be the time-evolved charge current. Then,
@@
$\begin{split}\mathcal{J}_{i}(t):=\alpha_{t}\Big{(}\mathrm{d}\Gamma(J_{i})\Big{)}\;,\qquad\mathcal{J}^{\mathrm{spin}}_{L;i}:=\mathrm{d}\Gamma\big{(}(J_{i}^{\mathrm{conv}})_{L}\big{)}\;;\end{split}$
@@
Note that we can extend $\mathcal{K}^{\beta,L}_{ij}$ to a $\beta$-periodic function on $\mathbb{R}$, since in fact by the KMS property we have $\mathcal{K}^{\beta,L}_{ij}(0)=\mathcal{K}^{\beta,L}_{ij}(\beta)$. Accordingly, for any $\beta>0$ and $\eta>0$ we let
@@
Under the same assumptions of Corollary A.13, we have
@@
Consider the first term in the Kubo formula (A.13). Noting that $J_{i}^{\mathrm{conv}}$, $P$ and $J_{j}(t)$ are in $\mathcal{P}(\mathcal{H})$ and uniformly bounded in $t$, we have that $t\mapsto\mathrm{e}^{\eta t}[J_{i}^{\mathrm{conv}},J_{j}(t)]P$ is Bochner integrable for fixed $\eta>0$ over $(-\infty,0]$. By Lemma 3.2(iv) and by Lemma A.8 Eq. (A.8) we have
$\begin{split}\int_{-\infty}^{0}\mathrm{e}^{\eta t}\tau\big{(}[J_{i}^{\mathrm{conv}},J_{j}(t)]P\big{)}\mathrm{d}t&=\tau\Bigg{(}\int_{-\infty}^{0}\mathrm{e}^{\eta t}[J_{i}^{\mathrm{conv}},J_{j}(t)]P\mathrm{d}t\Bigg{)}\\ &=\lim_{L\to\infty}\lim_{\beta\to\infty}|C_{L}|^{-1}\int_{-\infty}^{0}\mathrm{e}^{\eta t}\omega^{H,\mu}_{\beta}\big{(}[\mathcal{J}^{\mathrm{spin}}_{L;i},\mathcal{J}_{j}(t)]\big{)}\mathrm{d}t\;,\end{split}$ (A.16)
where we also swapped integral and expectation with the grand canonical state because the operator is Bochner integrable. Since $[\mathcal{J}^{\mathrm{spin}}_{L;i},\mathcal{J}_{j}(t)]$ is uniformly bounded in $t$, for $\beta$ large enough ($\beta^{-1}\lesssim\eta$) we have
$\begin{split}\Big{|}\int_{-\infty}^{0}&\big{(}\mathrm{e}^{\eta t}-\mathrm{e}^{\eta_{\beta}t}\big{)}\omega^{H,\mu}_{\beta}\big{(}[\mathcal{J}^{\mathrm{spin}}_{L;i},\mathcal{J}_{j}(t)]\big{)}\mathrm{d}t\Big{|}\\ &\lesssim\int_{-\infty}^{0}\big{(}\mathrm{e}^{\eta t}-\mathrm{e}^{\eta_{\beta}t}\big{)}\mathrm{d}t\leq\eta^{-1}-\eta_{\beta}^{-1}\lesssim\frac{1}{\eta^{2}\beta}\;,\end{split}$ (A.17)
@@
$\begin{split}\int_{-\infty}^{0}\mathrm{e}^{\eta_{\beta}t}&\omega^{H,\mu}_{\beta}\Big{(}[\mathcal{J}^{\mathrm{spin}}_{L;i},\mathcal{J}_{j}(t)]\Big{)}\mathrm{d}t\\ &=\int_{-\infty}^{0}\mathrm{e}^{\eta_{\beta}t}\Big{(}K^{\beta,L}_{ij}(t)-K^{\beta,L}_{ij}(t+\mathrm{i}\beta)\Big{)}\mathrm{d}t\\ &=\lim_{T\to\infty}\int_{-T}^{0}\Big{(}\mathrm{e}^{\eta_{\beta}t}K^{\beta,L}_{ij}(t)-\mathrm{e}^{\eta_{\beta}(t+\mathrm{i}\beta)}K^{\beta,L}_{ij}(t+\mathrm{i}\beta)\Big{)}\\ &=\mathrm{i}\int_{0}^{\beta}\mathrm{e}^{\mathrm{i}\eta_{\beta}t}K_{ij}^{\beta,L}(\mathrm{i}t)-\mathrm{i}\lim_{T\to\infty}\mathrm{e}^{-\eta_{\beta}T}\int_{0}^{\beta}\mathrm{e}^{\mathrm{i}\eta_{\beta}t}K_{ij}^{\beta,L}(-T+\mathrm{i}t)\mathrm{d}t\end{split}$ (A.18)
where we used the Cauchy integral formula, $K^{\beta,L}_{ij}$ being analytic. Because $[\mathcal{J}^{\mathrm{spin}}_{L;i},\mathcal{J}_{j}(-T+\mathrm{i}t)]$ is uniformly bounded in $T\in\mathbb{R}$ and locally bounded in $t$, so is $K_{ij}^{\beta,L}$. Therefore, since $\eta_{\beta}>0$ by definition, the last term in (A.18) is identically zero. We also note that $\mathcal{J}^{\mathrm{spin}}_{L;i}\mathcal{J}_{j}(\mathrm{i}t)=\mathbf{T}\mathcal{J}^{\mathrm{spin}}_{L;i}\mathcal{J}_{j}(\mathrm{i}t)$ for $t>0$ and thus
@@
$\int_{-\infty}^{0}\mathrm{e}^{\eta t}\tau\big{(}[J_{i}^{\mathrm{conv}},J_{j}(t)]P\big{)}\mathrm{d}t=\mathrm{i}\lim_{L\to\infty}\lim_{\beta\to\infty}\frac{1}{|C_{L}|}\int_{0}^{\beta}\mathrm{e}^{\mathrm{i}\eta_{\beta}t}\mathcal{K}_{ij}^{\beta,L}(t)\;.$ (A.20)
Let us now analyse the second term in the Kubo formula (A.13). Let $\mathcal{X}_{j}:=\mathrm{d}\Gamma(X_{j})$. By Lemma A.8 we have
$\tau\big{(}[J_{i}^{\mathrm{conv}},X_{j}]P\big{)}=\lim_{L\to\infty}\lim_{\beta\to\infty}|C_{L}|^{-1}\omega^{H,\mu}_{\beta}\big{(}\big{[}\mathcal{J}^{\mathrm{spin}}_{L;i},\mathcal{X}_{j}\big{]}\big{)}$
We want to open the commutator and this can be done provided that we localize the operator $\mathcal{X}_{j}$ so that it is bounded. By the argument at the end of the proof of Lemma A.8 we note that for any $R\geq L$ we have $\omega^{H,\mu}_{\beta}\big{(}\big{[}\mathcal{J}^{\mathrm{spin}}_{L;i},\mathcal{X}_{j}\big{]}\big{)}=\omega^{H,\mu}_{\beta}\big{(}\big{[}\mathcal{J}^{\mathrm{spin}}_{L;i},\mathcal{X}_{R;j}\big{]}\big{)}$, where $\mathcal{X}_{R,j}:=\mathrm{d}\Gamma(\chi_{R}X_{j})$ is bounded when $R$ is finite. Then, by the KMS property we can write
@@
and, by noting that $\mathcal{J}_{R,j}(\mathrm{i}t)=-\mathrm{i}\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{X}_{R,j}(\mathrm{i}t)$, we can write
$\omega^{H,\mu}_{\beta}\big{(}\big{[}\mathcal{J}^{\mathrm{spin}}_{L;i},\mathcal{X}_{j}\big{]}\big{)}=-\mathrm{i}\,\omega^{H,\mu}_{\beta}\Big{(}\int_{0}^{\beta}\mathcal{J}^{\mathrm{spin}}_{L;i}\mathcal{J}_{R,j}(\mathrm{i}t)\mathrm{d}t\Big{)}=-\mathrm{i}\,\int_{0}^{\beta}\omega^{H,\mu}_{\beta}\big{(}\mathcal{J}^{\mathrm{spin}}_{L;i}\mathcal{J}_{R,j}(\mathrm{i}t)\big{)}\mathrm{d}t\;,$
where in the last step we swapped integration with expectation over the grand canonical state because the operator $\mathcal{J}_{L,i}(-\mathrm{i}t)\mathcal{J}_{R,j}$ is bounded, locally in $t$, and thus Bochner integrable over $[0,\beta]$. We shall now remove the localisation $R$ by taking the limit $R\to\infty$. This can be done by noting that $\mathcal{J}_{L,i}(-\mathrm{i}t)\mathcal{J}_{R,j}$ converges in the operator topology to $\mathcal{J}_{L,i}(-\mathrm{i}t)\mathcal{J}_{j}$, which is Bochner integrable over $[0,\beta]$. Thus, we can take the limit inside the integral by the Lebesgue dominated convergence theorem as well as inside $\omega^{H,\mu}_{\beta}$. Therefore, using (A.19) we can write
@@
To conclude, we obtain an explicit expression for $\widehat{\mathcal{K}}_{ij}$ and thus prove Theorem 4.15.
@@
Note that, while $H$ is invariant under $R_{v\otimes h}$, both $J^{\mathrm{conv}}_{1}$ and $J_{2}$ switch sign under such a symmetry; thus
$\omega^{H,\mu}_{\beta}(\mathcal{J}^{\mathrm{spin}}_{L,1})=\omega^{H,\mu}_{\beta}(\mathcal{J}_{2}(\mathrm{i}t))=0\;,$ (A.21)
and by invariance under rotations one can likewise prove that the other components $\mathcal{J}^{\mathrm{spin}}_{L,2}$ and $\mathcal{J}_{1}(\mathrm{i}t)$ have zero expectation. We introduce the finite-temperature imaginary-time-ordered Green function as the following family of operators on $\mathcal{H}$
$G^{H,\mu}_{\beta}(t):=\frac{\mathrm{e}^{-t(H-\mu\mathbb{1})}}{\mathbb{1}+\mathrm{e}^{-\beta(H-\mu\mathbb{1})}}\qquad t\in[0,\beta)\;,$
and extend it to the whole $t\in\mathbb{R}$ by the anti-periodicity condition $G^{H,\mu}_{\beta}(t-\beta)=-G^{H,\mu}_{\beta}(t)$. Note that since $\alpha_{t}(a^{*}(f))=a^{*}(\mathrm{e}^{\mathrm{i}t(H-\mu\mathbb{1})}f)$, by (A.1) we have for any $f,g\in\mathcal{H}$
$\omega^{H,\mu}_{\beta}\big{(}a(f)\alpha_{t}(a^{*}(g))\big{)}=\langle f,\big{(}\mathbb{1}+\mathrm{e}^{-\beta(H-\mu\mathbb{1})}\big{)}^{-1}\mathrm{e}^{\mathrm{i}t(H-\mu\mathbb{1})}g\rangle_{\mathcal{H}}\;.$
Therefore, thanks to the anti-periodicity condition, for any $f,g\in\mathcal{H}$ we have
$\langle f,G^{H,\mu}_{\beta}(t)g\rangle_{\mathcal{H}}=\omega^{H,\mu}_{\beta}\big{(}\mathbf{T}a(f)\alpha_{\mathrm{i}t}(a^{*}(g))\big{)}\;,$ (A.22)
which explains why the operators $G^{H,\mu}_{\beta}$ are called finite-temperature imaginary-time-ordered Green function. Note that because $(H,\mu)$ is an insulator $G^{H,\mu}_{\beta}(t)\in\mathcal{P}_{0}(\mathcal{H})$ for any $t\in\mathbb{R}$ and any $\beta>0$. We also note that $G^{H,\mu}_{\beta}(t)$ is $2\beta$-periodic so that it is convenient to introduce Fourier components. Because $G^{H,\mu}_{\beta}(t-\beta)=-G^{H,\mu}_{\beta}(t)$, only odd frequencies are non-zero; thus, for $\omega\in\frac{2\pi}{\beta}(\mathbb{Z}+\frac{1}{2})$ we set (5) (5) (5)Note that our convention for the Fourier transform is opposite with respect to [GMP1, GMP2, MaPo2].
@@
Since the grand canonical state is quasi-free, that is, it satisfies the fermionic Wick rule, by using (A.21) and (A.22), we find
$\begin{split}&\omega^{H,\mu}_{\beta}\big{(}\mathcal{J}^{\mathrm{spin}}_{L;i}\mathcal{J}_{j}(\mathrm{i}t)\big{)}=\operatorname{Tr}_{\mathcal{H}}\Big{(}\big{(}J_{i}^{\mathrm{conv}}\big{)}_{L}G^{H,\mu}_{\beta}(t)J_{j}G^{H,\mu}_{\beta}(-t)\Big{)}\;.\end{split}$
@@
$\widehat{\mathcal{K}}_{ij}(\eta)=(2\pi)^{-1}\lim_{L\to\infty}|C_{L}|^{-1}\int_{\mathbb{R}}\operatorname{Tr}_{\mathcal{H}}\Big{(}\big{(}J_{i}^{\mathrm{conv}}\big{)}_{L}\widehat{G}^{H,\mu}(\omega-\eta)J_{j}\widehat{G}^{H,\mu}(\omega)\Big{)}\;.$
@@
where $\mathbf{k}\equiv(\omega,k)$ and where $\widehat{G}^{H,\mu}(\omega;k):=\big{(}H(k)-(\mu-\mathrm{i}\omega)\mathbb{1}\big{)}^{-1}$. To conclude, since $\partial_{k_{i}}H(k)=\partial_{k_{i}}\widehat{G}^{H,\mu}(\omega;k)^{-1}$, it is easy to check that

### v2

@@ 4.3. Lack of quantisation
Note that for $m\neq 0$ the integrand in (4.25) is continuous in $\lambda_{\mathrm{R}}$, in a neighbourhood of 0, and uniformly bounded by some integrable function (decaying in $k_{0}$ and constant in $k_{1}$ and $k_{2}$). Thus, by the dominated convergence theorem $\sigma^{s}_{12}$ is also continuous in $\lambda_{\mathrm{R}}$, in a neighbourhood of 0.
@@ A.1. Grand canonical formalism
$\{a(f),a(g)\}=\{a^{*}(f),a^{*}(g)\}=0\;,\qquad\{a(f),a^{*}(g)\}=\langle f,g\rangle_{\mathcal{H}}\;,\qquad\forall f,g\in\mathcal{H}\;,$
@@
$\alpha_{t}\big(a(f)\big):=a(\mathrm{e}^{\mathrm{i}t(H-\mu\mathbb{1})}f)\;,\qquad\forall f\in\mathcal{H}\;,$ (A.3)
see [BR, Section 5.2.4]. More precisely, one introduces the set of states that satisfy the KMS property with respect to $(\alpha_{t})_{t\in\mathbb{R}}$, see [BR, Definition 5.3.1].
@@
Following the proof of Lemma A.10, one can see that the second and third terms are $O(L^{2-1/\xi})$, $\xi$ being the constant associated with $B$ according to (2.5), so that the first claim is proven. Concerning (A.10), since by Lemma 3.5 $B,X]\in\mathcal{P}(\mathcal{H})$ we can apply Lemma A.10 and use that $\chi_{L}$ commutes with $X$. ∎
@@
$\mathrm{d}\Gamma(A_{L}),\mathrm{d}\Gamma(B)]=\mathrm{d}\Gamma([A_{L},B])=\lim_{R\to\infty}\mathrm{d}\Gamma([A_{L},B_{R}])$
@@
since $\gamma^{H,\mu}_{\beta}$ is bounded. Then, we can take the $\beta\to\infty$ as in (A.12) since $A_{L},B]$ is trace class, and conclude by (A.9).
Finally, to prove the last claim we use that $\chi_{L}$ is a projection and that it commutes with $X$, so that $B_{L},X]=[B_{L},X_{L}]$. But $X_{L}$ is bounded hence by (A.6)
@@
where we also used that $X=X_{L}+\chi^{c}_{L}X$ and that $\mathrm{d}\Gamma(B_{L}),\mathrm{d}\Gamma(\chi^{c}_{L}X)]=\mathrm{d}\Gamma\big([B_{L},\chi^{c}_{L}X]\big)=0$. The conclusion follows by arguing as before, and by using (A.10). ∎
@@ A.2. Wick’s rotation of the Kubo formula
To begin, we provide a Kubo formula for the spin conductivity in $\mathcal{P}_{0}^{\hexagon}(\mathcal{H})$. We henceforth denote $J:=\mathrm{i}[H,X]$ the charge current operator associated with the Hamiltonian $H$.
@@
Let $(H,\mu)$ be an insulator with $H\in\mathcal{P}_{0}^{\hexagon}(\mathcal{H})$ and let $P=P(H,\mu)$ be the associated Fermi projector. Let $J_{i}^{\mathrm{conv}}$ be the conventional spin current and let $J(s):=\mathrm{e}^{\mathrm{i}sH}J\mathrm{e}^{-\mathrm{i}sH}$ be the time-evolved charge current. Then,
@@
[…]
@@
For any $\beta>0$ and $\eta\geq 0$ we let
@@
Under the same assumptions of Proposition A.13, we have
@@
Consider the first term in the Kubo formula (A.13). By Lemma 3.4 and Lemma 3.5, we note that $J_{i}^{\mathrm{conv}}$, $P$ and $J_{j}(t)$ are in $\mathcal{P}_{0}(\mathcal{H})$ for any $t\in\mathbb{R}$. It is then easy to check that $\mathcal{J}^{\mathrm{spin}}_{L;i},\mathcal{J}_{j}(t)]$ is uniformly (w.r.t. $t\in\mathbb{R}$) bounded as an operator on $\mathcal{F}(\mathcal{H})$, as long as $L$ is finite. Thus, by using the first identity in (A.8) we have
$\begin{split}\int_{-\infty}^{0}\mathrm{e}^{\eta t}\tau\big([J_{i}^{\mathrm{conv}},J_{j}(t)]P\big)\mathrm{d}t&=\int_{-\infty}^{0}\mathrm{e}^{\eta t}\lim_{L\to\infty}\lim_{\beta\to\infty}|C_{L}|^{-1}\omega^{H,\mu}_{\beta}\big([\mathcal{J}^{\mathrm{spin}}_{L;i},\mathcal{J}_{j}(t)]\big)\mathrm{d}t\\ &=\lim_{L\to\infty}\lim_{\beta\to\infty}|C_{L}|^{-1}\int_{-\infty}^{0}\mathrm{e}^{\eta t}\omega^{H,\mu}_{\beta}\big([\mathcal{J}^{\mathrm{spin}}_{L;i},\mathcal{J}_{j}(t)]\big)\mathrm{d}t\;,\end{split}$ (A.16)
where we swapped the limit with the integral by the dominated convergence theorem. Using once more that $\mathcal{J}^{\mathrm{spin}}_{L;i},\mathcal{J}_{j}(t)]$ is uniformly bounded in $t$, we have
$\begin{split}\Big|\int_{-\infty}^{0}&\big(\mathrm{e}^{\eta t}-\mathrm{e}^{\eta_{\beta}t}\big)\omega^{H,\mu}_{\beta}\big([\mathcal{J}^{\mathrm{spin}}_{L;i},\mathcal{J}_{j}(t)]\big)\mathrm{d}t\Big|\\ &\lesssim\int_{-\infty}^{0}\big(\mathrm{e}^{\eta t}-\mathrm{e}^{\eta_{\beta}t}\big)\mathrm{d}t=\eta^{-1}-\eta_{\beta}^{-1}\leq\frac{2\pi}{\eta^{2}\beta}\;,\end{split}$ (A.17)
@@
$\begin{split}\int_{-\infty}^{0}\mathrm{e}^{\eta_{\beta}t}&\omega^{H,\mu}_{\beta}\Big([\mathcal{J}^{\mathrm{spin}}_{L;i},\mathcal{J}_{j}(t)]\Big)\mathrm{d}t\\ &=\int_{-\infty}^{0}\mathrm{e}^{\eta_{\beta}t}\Big(K^{\beta,L}_{ij}(t)-K^{\beta,L}_{ij}(t+\mathrm{i}\beta)\Big)\mathrm{d}t\\ &=\lim_{T\to\infty}\int_{-T}^{0}\Big(\mathrm{e}^{\eta_{\beta}t}K^{\beta,L}_{ij}(t)-\mathrm{e}^{\eta_{\beta}(t+\mathrm{i}\beta)}K^{\beta,L}_{ij}(t+\mathrm{i}\beta)\Big)\mathrm{d}t\\ &=\mathrm{i}\int_{0}^{\beta}\mathrm{e}^{\mathrm{i}\eta_{\beta}t}K_{ij}^{\beta,L}(\mathrm{i}t)\mathrm{d}t-\mathrm{i}\lim_{T\to\infty}\mathrm{e}^{-\eta_{\beta}T}\int_{0}^{\beta}\mathrm{e}^{\mathrm{i}\eta_{\beta}t}K_{ij}^{\beta,L}(-T+\mathrm{i}t)\mathrm{d}t\end{split}$ (A.18)
where we used the Cauchy integral formula for $K^{\beta,L}_{ij}$ being analytic. Because $\mathcal{J}^{\mathrm{spin}}_{L;i},\mathcal{J}_{j}(-T+\mathrm{i}t)]$ is uniformly bounded in $T\in\mathbb{R}$ and locally bounded in $t$, so is $K_{ij}^{\beta,L}$. Therefore, since $\eta_{\beta}>0$ by definition, the last term in (A.18) is identically zero. We also note that $\mathcal{J}^{\mathrm{spin}}_{L;i}\mathcal{J}_{j}(\mathrm{i}t)=\mathbf{T}\mathcal{J}^{\mathrm{spin}}_{L;i}\mathcal{J}_{j}(\mathrm{i}t)$ for $t>0$ and thus
@@
$\int_{-\infty}^{0}\mathrm{e}^{\eta t}\tau\big([J_{i}^{\mathrm{conv}},J_{j}(t)]P\big)\mathrm{d}t=\mathrm{i}\lim_{L\to\infty}\lim_{\beta\to\infty}\frac{1}{|C_{L}|}\int_{0}^{\beta}\mathrm{e}^{\mathrm{i}\eta_{\beta}t}\mathcal{K}_{ij}^{\beta,L}(t)\mathrm{d}t\;.$ (A.20)
Let us now analyse the second term in the Kubo formula (A.13). Let $\mathcal{X}_{j}:=\mathrm{d}\Gamma(X_{j})$. By employing the second identity in (A.8), we have
$\tau\big([J_{i}^{\mathrm{conv}},X_{j}]P\big)=\lim_{L\to\infty}\lim_{\beta\to\infty}|C_{L}|^{-1}\omega^{H,\mu}_{\beta}\big(\big[\mathcal{J}^{\mathrm{spin}}_{L;i},\mathcal{X}_{j}\big]\big).$
We want to open the commutator on the right-hand side term in the last equality and this can be done provided that we localize the operator $\mathcal{X}_{j}$ so that it is bounded. By the argument at the end of the proof of Lemma A.8 we note that for any $R\geq L$ we have $\omega^{H,\mu}_{\beta}\big(\big[\mathcal{J}^{\mathrm{spin}}_{L;i},\mathcal{X}_{j}\big]\big)=\omega^{H,\mu}_{\beta}\big(\big[\mathcal{J}^{\mathrm{spin}}_{L;i},\mathcal{X}_{R;j}\big]\big)$, where $\mathcal{X}_{R;j}:=\mathrm{d}\Gamma(\chi_{R}X_{j})$ is bounded when $R$ is finite. Then, by the KMS property we can write
@@
and, by slight abuse of notation, letting $\mathcal{J}_{R;j}(\mathrm{i}t):=-\mathrm{i}\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{X}_{R;j}(\mathrm{i}t)=\alpha_{\mathrm{i}t}\big(\mathrm{d}\Gamma(\mathrm{i}[H,\chi_{R}X_{j}])\big)$, we can write
$\omega^{H,\mu}_{\beta}\big(\big[\mathcal{J}^{\mathrm{spin}}_{L;i},\mathcal{X}_{j}\big]\big)=-\mathrm{i}\,\omega^{H,\mu}_{\beta}\Big(\int_{0}^{\beta}\mathcal{J}^{\mathrm{spin}}_{L;i}\mathcal{J}_{R;j}(\mathrm{i}t)\mathrm{d}t\Big)=-\mathrm{i}\,\int_{0}^{\beta}\omega^{H,\mu}_{\beta}\big(\mathcal{J}^{\mathrm{spin}}_{L;i}\mathcal{J}_{R;j}(\mathrm{i}t)\big)\mathrm{d}t\;,$
where in the last step we swapped integration with expectation over the grand canonical state because the operator $\mathcal{J}^{\mathrm{spin}}_{L;i}\mathcal{J}_{R;j}(\mathrm{i}t)$ is bounded, locally in $t$, and thus Bochner integrable over $0,\beta]$. We shall now remove the localisation $R$ by taking the limit $R\to\infty$. This can be done by noting that $\mathcal{J}_{L;i}\mathcal{J}_{R;j}(\mathrm{i}t)$ converges in the operator topology to $\mathcal{J}_{L;i}\mathcal{J}_{j}(\mathrm{i}t)$, which is Bochner integrable over $0,\beta]$. Thus, we can take the limit inside the integral by the Lebesgue dominated convergence theorem as well as inside $\omega^{H,\mu}_{\beta}$. Therefore, using (A.19) we can write
@@
To conclude, we obtain an explicit expression for $\widehat{\mathcal{K}}_{ij}$ and thus we can prove Theorem 4.15.
@@
We compute $\mathcal{K}^{\beta,L}_{ij}(t)=\omega^{H,\mu}_{\beta}\big(\mathbf{T}\mathcal{J}^{\mathrm{spin}}_{L;i}\mathcal{J}_{j}(\mathrm{i}t)\big)=\omega^{H,\mu}_{\beta}\big(\mathcal{J}^{\mathrm{spin}}_{L;i}\mathcal{J}_{j}(\mathrm{i}t)\big)$, $t\in(0,\beta]$, by using that $\omega^{H,\mu}_{\beta}$ is quasi-free. First of all, we note that
$\omega^{H,\mu}_{\beta}(\mathcal{J}_{j}(\mathrm{i}t))=0\;,\qquad j=1,2\;.$ (A.21)
This follows by observing that $H$ is invariant under $2\pi/3$-rotations, whereas $J_{i}$ transform like vectors. By the Wick rule and simple algebraic manipulations (7) (7) (7)One can indeed check that the Wick rule holds for the expectations with respect to $\omega^{H,\mu}_{\beta}$ of time-ordered products in the creation and annihilation operators., for $t\in(0,\beta]$ we find
$\begin{split}&\omega^{H,\mu}_{\beta}\big(\mathbf{T}\mathcal{J}^{\mathrm{spin}}_{L;i}\mathcal{J}_{j}(\mathrm{i}t)\big)=\operatorname{Tr}_{\mathcal{H}}\Big(\big(J_{i}^{\mathrm{conv}}\big)_{L}G^{H,\mu}_{\beta}(t)J_{j}G^{H,\mu}_{\beta}(-t)\Big)\;.\end{split}$
where the operator $G^{H,\mu}_{\beta}(t)$ is the finite-temperature imaginary-time-ordered Green function, that is, for any $f,g\in\mathcal{H}$
$\langle f,G^{H,\mu}_{\beta}(t)g\rangle_{\mathcal{H}}:=\omega^{H,\mu}_{\beta}\big(\mathbf{T}a(f)\alpha_{\mathrm{i}t}(a^{*}(g))\big)\;.$ (A.22)
By (A.1) and by noting that $\alpha_{\mathrm{i}t}(a^{*}(f))=a^{*}(\mathrm{e}^{-t(H-\mu\mathbb{1})}f)$, we find that $G^{H,\mu}_{\beta}(t)$ is given by
$G^{H,\mu}_{\beta}(t):=\frac{\mathrm{e}^{-t(H-\mu\mathbb{1})}}{\mathbb{1}+\mathrm{e}^{-\beta(H-\mu\mathbb{1})}},\qquad t\in[0,\beta)\;,$
while in the rest of the interval $\beta,\beta]$ it is determined by the anti-periodicity condition $G^{H,\mu}_{\beta}(t-\beta)=-G^{H,\mu}_{\beta}(t)$. Note that, by Lemma 3.5, $G^{H,\mu}_{\beta}(t)\in\mathcal{P}_{0}(\mathcal{H})$ for any $t\in\mathbb{R}$ and any $\beta>0$. We also introduce the Fourier components of $G^{H,\mu}_{\beta}(t)$; because $G^{H,\mu}_{\beta}(t-\beta)=-G^{H,\mu}_{\beta}(t)$, only odd frequencies are non-zero; thus, for $\omega\in\frac{2\pi}{\beta}(\mathbb{Z}+\frac{1}{2})$ we set (8) (8) (8)Note that our convention for the Fourier transform is opposite with respect to [GMP1, GMP2, MaPo2].
@@
$\widehat{\mathcal{K}}_{ij}(\eta)=(2\pi)^{-1}\lim_{L\to\infty}|C_{L}|^{-1}\int_{\mathbb{R}}\mathrm{d}\omega\operatorname{Tr}_{\mathcal{H}}\Big(\big(J_{i}^{\mathrm{conv}}\big)_{L}\widehat{G}^{H,\mu}(\omega-\eta)J_{j}\widehat{G}^{H,\mu}(\omega)\Big)\;.$
@@
where $\mathbf{k}\equiv(\omega,k)$ and where $\widehat{G}^{H,\mu}(\omega;k):=\big(H(k)-(\mu-\mathrm{i}\omega)\mathbb{1}\big)^{-1}$. To conclude, since $\partial_{k_{i}}H(k)=\partial_{k_{i}}\left(\widehat{G}^{H,\mu}(\omega;k)^{-1}\right)$, it is easy to check that